# scale cd/.style=every label/.append style=scale=#1, cells=nodes=scale=#1

### v1

@@ 4. Block-theoretic preliminaries
Throughout, let $G$ and $H$ be finite groups and assume $(K,\mathcal{O},k)$ is large enough for $G\times H$, with $k$ algebraically closed. We review some important preliminaries regarding Brauer pairs and block fusion systems. We assume the reader is familiar with the definition of a block, and that blocks of $\mathcal{O}G$ and $kG$ are in one-to-one correspondence induced by the canonical surjection. If $B$ is a sum of blocks of $kG$, then we write $e_{B}$ for the identity element of $B$.
@@
We first state some well-known structural results, see [10] for details.
@@
This is [4, Proposition 4.3], with the unused assumption that $B$ is a block algebra removed. ∎
@@
We next recall the definition of a fusion system of a block. See [2] for more details, proofs, and the definition of an abstract fusion system and a saturated fusion system. Let $B$ be a block of $kG$ and let $(P,e)$ be a maximal $B$-Brauer pair. For $Q\leq P$, denote by $e_{Q}$ the unique block idempotent of $kC_{G}(Q)$ for which $(Q,e_{Q})\leq(P,e)$. The fusion system of $B$ associated to $(P,e)$ is the category $\mathcal{F}$ whose objects are subgroups of $P$ and whose morphism set $\operatorname{Hom}_{\mathcal{F}}(Q,R)$ is the set of group homomorphisms arising as conjugation maps $c_{g}:Q\to R$ where $g\in G$ satisfies (${}^{g}(Q,e_{Q})\leq(R,e_{R})$. It is well known that $\mathcal{F}$ is a saturated fusion system on $P$. For an abstract fusion system $\mathcal{F}$ on a $p$-group $P$, a subgroup $Q\leq P$ is fully $\mathcal{F}$-centralized (resp. fully $\mathcal{F}$-normalized) if $|C_{P}(Q)|\geq|C_{P}(Q^{\prime})|$ for all (resp. $|N_{P}(Q)|\geq N_{P}(Q^{\prime})$) for all subgroups $Q^{\prime}$ of $P$ which are $\mathcal{F}$-isomorphic to $Q$. Moreover, $Q$ is $\mathcal{F}$-centric if $C_{P}(Q^{\prime})=Z(Q^{\prime})$ for all $Q^{\prime}$ that are $\mathcal{F}$-isomorphic to $Q$.
The following theorem of Alperin and Brouè in [1] characterizes $\mathcal{F}$-centric, fully normalized, and fully centralized subgroups for block fusion systems. We give the formulation presented in [12].
@@
[12, Theorems 2.4, 2.5] Let $B$ be a block of $kG$, let $(P,e)$ be a maximal $B$-Brauer pair, and let $\mathcal{F}$ be the fusion system associated to $B$ and $(P,e)$. For every subgroup $Q\leq P$, denote by $e_{Q}$ the unique block idempotent of $k[C_{G}(Q)]$ for which $(Q,e_{Q})\leq(P,e)$.
@@ 5. Trivial source rings, $p$-permutation equivalences, and splendid Rickard equivalences
    We denote by $T(RGe)$ the trivial source ring, which is the Grothendieck group of the category ${}_{RGe}\textbf{triv}$ with respect to direct sums. The group $T(RGe)$ is free as an abelian group with standard basis given by the elements $[M]$, where $M$ runs through a set of representatives of isomorphism classes of trivial source $RG$-modules. It forms a commutative ring under multiplication induced by $\otimes_{R}$, denoted $\cdot_{R}$. We always view $T(RGe)$ as a subgroup of $T(RG)$ in the natural way. Moreover, multiplying a $p$-permutation $RG$-module by $e$ induces a projection $T(RG)\to T(RGe)$. In this way, if $b_{1},\dots,b_{k}$ are the blocks of $RG$, then we have a direct sum decomposition of abelian groups $T(RG)=T(RGb_{1})\oplus\cdots\oplus T(RGb_{k})$. Taking duals induces a self-inverse group isomorphism $T(RG)\to T(RG)$. We write $O(T(RG))$ for the orthogonal unit group of the trivial source ring, that is, the subgroup of $T(RG)^{\times}$ consisting of all virtual $p$-permutation modules $\gamma$ whose inverse is its dual, i.e. $\gamma\cdot_{R}\gamma^{*}=[R]=\gamma^{*}\cdot_{R}\gamma$. This notion does not extend to idempotents because if $\gamma\in O(T(RGe))$, then $\gamma^{*}\in O(T(RGe^{*}))$, where $(-)^{*}$ is the map induced by the involution $g\mapsto g^{-1}$.
@@
    Similarly, we define $T(RGe,RHf):=T(R[G\times H])(e\otimes f^{*})$ where we identify $R[G\times H]$ with $RG\otimes_{R}RH$ in the natural way. In the case where $e=f$, $T(RGe,RGe)$ has a ring structure induced by the tensor product $\otimes_{RGe}$, written $\cdot_{G}$. We denote by $T^{\Delta}(RGe,RHf)$ the subgroup of $T(RGe,RHf)$ the subgroup generated by all trivial source $(RGe,RHf)$-bimodules which have twisted diagonal vertices. Taking bimodule duals induces a group isomorphism $T(RGe,RHf)\to T(RHf,RGe)$, which has inverse again induced by taking duals. Denote by $O(T(RGe,RGe))$ the subgroup of $T(RGe,RGe)^{\times}$ consisting of all virtual $p$-permutation bimodules $\gamma$ whose inverse is its dual, i.e. $\gamma\cdot_{G}\gamma^{*}=[RGe]=\gamma^{*}\cdot_{G}\gamma$. Finally, observe that any $M\in T(RGe,RHf)$ induces a group homomorphism $M\cdot_{H}-:T(RHf)\to T(RGe)$. If $M\in T(RGe,RHf)^{\times}$, it is an isomorphism.
@@
    The following definition is due to Boltje and Xu in [5]. A $p$-permutation equivalence between $RGe$ and $RHf$ is an element $\gamma\in T^{\Delta}(RGe,RHf)$ satisfying
@@
Let $A$ be a sum of blocks of $RG$ and $B$ a sum of blocks of $RH$. A splendid Rickard complex for $A$ and $B$ is a bounded chain complex $C$ of $p$-permutation $(A,B)$-bimodules which satisfy the following conditions:
@@
    $C\otimes_{B}C^{*}\simeq A$ in $K^{b}({}_{A}\mathbf{mod}_{A})$ and $C^{*}\otimes_{A}C\simeq B$ in $K^{b}({}_{B}\mathbf{mod}_{B})$, where $A$ and $B$ are considered as chain complexes concentrated in degree 00.
A splendid Rickard complex $C$ of $(A,B)$-bimodules induces an equivalence of categories $C\otimes_{B}-:K^{b}({}_{B}\textbf{mod})\to K^{b}({}_{A}\textbf{mod})$. We call this resulting equivalence a splendid equivalence between $A$ and $B$.
@@
Since $B$ is symmetric and the terms of $C$ are projective as right $B$-modules, we have an isomorphism $C\otimes_{B}C^{*}\cong\mathrm{End}_{B^{op}}(C)$ of chain complexes of $(A,A)$-bimodules. Therefore, $\mathrm{End}_{B^{op}}(C)$ is contractible by the hypothesis. The complex $\mathrm{End}_{A\otimes_{k}B^{op}}(C)$ is obtained from the complex $\mathrm{End}_{B^{op}}(C)$ by taking $A$-fixed points. Since taking $A$-fixed points is an additive functor, $\mathrm{End}_{A\otimes_{R}B^{op}}(C)$ is also contractible. Taking homology we obtain
@@
If $A$ and $B$ are blocks of $RG$ and $RH$, then a splendid Rickard complex for $A$ and $B$ is an indecomposable object in the homotopy category of chain complexes of $(A,B)$-bimodules.
@@
Suppose that $A$ is a block of $RG$ and $B$ is a block of $RH$. Let $C$ be a splendid Rickard complex for $A$ and $B$. Then $C$ is indecomposable in $K^{b}({}_{A}\mathbf{mod}_{B})$. In particular, $C$ has, up to isomorphism, a unique indecomposable noncontractible direct summand which is a splendid Rickard equivalence for $A$ and $B$.
@@
The notion of a “splendid” (split endomorphism two-sided tilting complex induced from diagonal subgroups) Rickard complex is due to Jeremy Rickard. In Rickard’s paper [16] such complexes are referred to as splendid equivalences, however the nomenclature “Rickard complex” had been used even prior to the publishing of [16], see for instance [9]. Rickard’s original definition assumed that $G$ and $H$ shared a common Sylow $p$-subgroup, and that the vertices of the modules in a splendid Rickard complex are contained inside the diagonal subgroup generated by the common Sylow. We remove that assumption here, instead asserting only that the modules are twisted diagonal.
@@
Assume that $A$ and $B$ are blocks of $RG$ and $RH$, respectively. Let $C$ be a bounded chain complex of $p$-permutation $(A,B)$-bimodules which satisfies condition 5.2(a). If $C\otimes_{B}C^{*}\simeq A$ in $K^{b}({}_{A}\mathbf{mod}_{A})$, then $C$ is a splendid Rickard complex.
@@
An indecomposable $(A,B)$-bimodule $M$ with a twisted diagonal vertex $X$ is projective as a left $A$-module and a right $B$-module. Indeed, by the Mackey Formula and the Krull-Schmidt theorem, every indecomposable direct summand of $\mathrm{Res}_{G\times 1}^{G\times H}(M)$ is projective relative to $(G\times 1)\cap{}^{(g,h)}X$ for some $(g,h)\in G\times H$. Since ${}^{(g,h)}X$ is also twisted diagonal, this intersection is trivial. This shows that $M$ is projective as a left $A$-module. A similar argument shows that $M$ is projective as a right $B$-module. Hence, the result follows from [13, 4.14.16]. ∎
@@
[5, Theorem 1.5] Let $C$ be a splendid Rickard complex between $A$ and $B$. Then $\gamma:=\Lambda(C)$ is a $p$-permutation equivalence between $RGe$ and $RHf$.
@@
Under Rickard’s original definition, if $C$ is a splendid Rickard complex for $A$ and $B$, for any $Q\leq P$ and blocks $e_{Q}$ and $f_{Q}$ corresponding to the blocks of $kC_{G}(Q)$ and $kC_{H}(Q)$ determined by $A$ and $B$, the chain complex $e_{Q}C(\Delta Q)f_{Q}$ is a splendid Rickard complex for $kC_{G}(Q)e_{Q}$ and $kC_{H}(Q)f_{Q}$. Such equivalences are regarded as “local” equivalences, equivalences between blocks of local centralizer subgroups.
@@ 6. Extended tensor products and tensor products of $p$-permutation bimodules
where $h\in H$ is chosen such that $(g,h)\in X$ and $(h,k)\in Y$. Note this is independent of choice of $h$. This construction was first used in [7]. We denote this extended tensor product by $M\underset{RH}{\overset{X,Y}{\otimes}}N\in{}_{R[X\ast Y]}\mathbf{mod}$, and obtain a bifunctor $-\underset{RH}{\overset{X,Y}{\otimes}}-:{}_{RX}\mathbf{mod}\times{}_{RY}\mathbf{mod}\to{}_{R[X\ast Y]}\mathbf{mod}$. Some quick calculations determine that this construction is associative in the following way: if $L$ is a finite group, $Z\leq K\times L$, and $P\in{}_{RZ}\mathbf{mod}$, then
@@
defines a natural isomorphism of $R[X\ast Y\ast Z]$-modules which is natural in $M,N$ and $K$. Moreover, for $g\in G$, $h\in H$, $k\in K$, we have a natural isomorphism
@@
[7] Let $X\leq G\times H$, $Y\leq H\times K$, $M\in{}_{RX}\mathbf{mod}$, $N\in{}_{RY}\mathbf{mod}$. There is a natural isomorphism
@@
[4, Lemma 7.2] Let $X\leq G\times H$ and $Y\leq H\times K$.
@@
This construction originally appears in [3]. For a fixed twisted diagonal $p$-subgroup $\Delta(P,\sigma,R)\leq G\times K$, denote by $\Gamma_{H}(P,\sigma,R)$ the set of triples $(\phi,Q,\psi)$ where $Q\leq H$ and $\psi:R\xrightarrow{\sim}Q$ and $\phi:Q\xrightarrow{\sim}P$ are isomorphisms with $\phi\circ\psi=\sigma$. The group $H$ acts on $\Gamma_{H}(P,\sigma,R)$ by
(${}^{h}(\phi,Q,\psi):=(\phi c_{h}^{-1},{}^{h}Q,c_{h}\psi).$
@@
(see [3, 3.1]) which is natural in $M$ and $N$. The following theorem characterizes the tensor product on the right-hand side for a particular class of $p$-permutation modules.
@@
[3, Theorem 3.3] Let $\Delta(P,\sigma,R)\in s_{p}(G\times K)$ be a twisted diagonal $p$-subgroup, let $\Gamma=\Gamma_{H}(P,\sigma,R)$ be as above, and let $\widetilde{\Gamma}\subseteq\Gamma$ be a set of representatives of the $H$-orbits of $\Gamma$. Furthermore, let $M\in{}_{\mathcal{O}G}\mathbf{triv}_{\mathcal{O}H}$ and $N\in{}_{\mathcal{O}H}\mathbf{triv}_{\mathcal{O}K}$ be $p$-permutation bimodules whose indecomposable direct summands have twisted diagonal vertices. Then the direct sum of the homomorphisms $\Phi_{(\phi,Q,\psi)}$ for $(\phi,Q,\psi)\in\widetilde{\Gamma}$ yields an isomorphism
@@
[4, Corollary 7.4] Let $M\in{}_{\mathcal{O}G}\mathbf{triv}_{\mathcal{O}H},N\in{}_{\mathcal{O}H}\mathbf{triv}_{\mathcal{O}K}$, $\Delta(P,\sigma,R)\in s_{p}(G\times K)$, and $\widetilde{\Gamma}\subseteq\Gamma=\Gamma_{H}(P,\sigma,R)$ be as before.
@@
    For the induced action of $N_{G\times K}(\Delta(P,\sigma,R))$ on the $H$-orbits $[\phi,Q,\psi]_{H}$ of $\Gamma$, one has
@@
[4, Theorem 7.5] Let $(P,e)\in\mathcal{BP}(kG)$ and $(R,d)\in\mathcal{BP}(kK)$. Suppose $\sigma:R\xrightarrow{\sim}P$ is an isomorphism and let $C_{G}(P)\leq S\leq N_{G}(P,e)$ and $C_{K}(R)\leq T\leq N_{K}(R,d)$ be intermediate subgroups. Further, let $\Omega:=\Omega_{H}((P,e),\sigma,(R,d))$ denote the set of triples $(\phi,(Q,f),\psi)$, where $(Q,f)\in\mathcal{BP}(kH)$ and $\psi:R\xrightarrow{\sim}Q$ and $\phi:Q\xrightarrow{\sim}P$ are isomorphisms such that $\sigma=\phi\circ\psi$. Finally, let $M\in{}_{\mathcal{O}G}\mathbf{triv}_{\mathcal{O}H}$ and $N\in{}_{\mathcal{O}H}\mathbf{triv}_{\mathcal{O}K}$ be $p$-permutation modules all of whose indecomposable direct summands have twisted diagonal vertices.

### v2

@@ 4. Block-theoretic preliminaries
Throughout, let $G$ and $H$ be finite groups and assume the $p$-modular system $(K,\mathcal{O},k)$ is large enough for $G\times H$ with $k$ algebraically closed for simplicity. We review some important preliminaries regarding Brauer pairs and block fusion systems. We assume the reader is familiar with the definition of a block, and that blocks of $\mathcal{O}G$ and $kG$ are in one-to-one correspondence induced by the canonical surjection. If $B$ is a sum of blocks of $kG$, then we write $e_{B}$ for the identity element of $B$.
@@
    If $e$ is a central idempotent of $kG$, then $\operatorname{br}_{P}(e)$ is an $N_{G}(P)$-stable central idempotent of $kC_{G}(P)$ and also a central idempotent of $kN_{G}(P)$. If $B$ is a sum of blocks of $kG$, we say that a Brauer pair $(P,e)$ is a $B$-Brauer pair if $\operatorname{br}_{P}(e_{B})e=e$, or equivalently if $\operatorname{br}_{P}(e_{B})e\neq 0$. Every Brauer pair is a $B$-Brauer pair for a unique block $B$ of $kG$, and in this case, $(\{1\},e_{B})\leq(P,e)$. The set of $B$-Brauer pairs is closed under conjugation and if $(Q,f)\leq(P,e)$ are Brauer pairs, $(Q,f)$ is a $B$-Brauer pair if and only if $(P,e)$ is. The set of $B$-Brauer pairs is denoted by $\mathcal{B}\mathcal{P}(B)$.
    For a sum $B$ of blocks of $\mathcal{O}G$, we simply define $\mathcal{B}\mathcal{P}(B)=\mathcal{B}\mathcal{P}(\overline{B})$. Thus Brauer pairs are by default viewed as pairs $(P,e)$, where $e$ is an idempotent of a group algebra over $k$. It may sometimes be convenient to lift the idempotent to the corresponding idempotent over $\mathcal{O}$, and we denote the set of resulting pairs by $\mathcal{B}\mathcal{P}_{\mathcal{O}}(B)$.
@@
We first state some well-known structural results, see [BrP80] for details.
@@
This is [BoPe20, Proposition 4.3], with the unused assumption that $B$ is a block algebra removed. ∎
@@
We next recall the definition of a fusion system of a block. See [AKO11] for more details, proofs, and the definition of an abstract fusion system and a saturated fusion system. Let $B$ be a block of $kG$ and let $(P,e)$ be a maximal $B$-Brauer pair. For $Q\leq P$, denote by $e_{Q}$ the unique block idempotent of $kC_{G}(Q)$ for which $(Q,e_{Q})\leq(P,e)$.
The fusion system of $B$ associated to $(P,e)$ is the category $\mathcal{F}$ whose objects are subgroups of $P$ and whose morphism set $\operatorname{Hom}_{\mathcal{F}}(Q,R)$ is the set of group homomorphisms arising as conjugation maps $c_{g}:Q\to R$ where $g\in G$ satisfies $^{g}(Q,e_{Q})\leq(R,e_{R})$. It is well known that $\mathcal{F}$ is a saturated fusion system on $P$.
For an abstract fusion system $\mathcal{F}$ on a $p$-group $P$, a subgroup $Q\leq P$ is fully $\mathcal{F}$-centralized (resp. fully $\mathcal{F}$-normalized) if $|C_{P}(Q)|\geq|C_{P}(Q^{\prime})|$ for all (resp. $|N_{P}(Q)|\geq|N_{P}(Q^{\prime})|$) for all subgroups $Q^{\prime}$ of $P$ which are $\mathcal{F}$-isomorphic to $Q$. Moreover, $Q$ is $\mathcal{F}$-centric if $C_{P}(Q^{\prime})=Z(Q^{\prime})$ for all $Q^{\prime}$ that are $\mathcal{F}$-isomorphic to $Q$.
The following theorem of Alperin and Broué in [AB79] characterizes $\mathcal{F}$-centric, fully normalized, and fully centralized subgroups for block fusion systems. We give the formulation presented in [L06].
@@
[L06, Theorems 2.4, 2.5] Let $B$ be a block of $kG$, let $(P,e)$ be a maximal $B$-Brauer pair, and let $\mathcal{F}$ be the fusion system associated to $B$ and $(P,e)$. For every subgroup $Q\leq P$, denote by $e_{Q}$ the unique block idempotent of $k[C_{G}(Q)]$ for which $(Q,e_{Q})\leq(P,e)$.
@@ 5. Trivial source rings, $p$-permutation equivalences, and splendid Rickard equivalences
    We denote by $T(RGe)$ the trivial source group for $RGe$, the Grothendieck group of the category $_{RGe}\textbf{triv}$ with respect to direct sums. The group $T(RGe)$ is free as an abelian group with standard basis given by the elements $[M]$, where $M$ runs through a set of representatives of isomorphism classes of trivial source $RGe$-modules. If $e=1$, then $T(RG)$ forms a commutative ring under multiplication induced by $\otimes_{R}$, denoted $\cdot_{R}$; we call $T(RG)$ the trivial source ring. We always view $T(RGe)$ as a subgroup of $T(RG)$ in the natural way. Moreover, multiplying a $p$-permutation RGRG-module by $e$ induces a projection $T(RG)\to T(RGe)$. In this way, if $b_{1},\dots,b_{k}$ are the blocks of $RG$, then we have a direct sum decomposition of abelian groups $T(RG)=T(RGb_{1})\oplus\cdots\oplus T(RGb_{k})$.
    Taking duals induces a self-inverse group isomorphism $T(RG)\to T(RG)$. We write $O(T(RG))$ for the orthogonal unit group of the trivial source ring, that is, the subgroup of $T(RG)^{\times}$ consisting of all virtual $p$-permutation modules $\gamma$ whose inverse is its dual, i.e. $\gamma\cdot_{R}\gamma^{*}=[R]=\gamma^{*}\cdot_{R}\gamma$. This notion does not extend to idempotents because if $\gamma\in O(T(RGe))$, then $\gamma^{*}\in O(T(RGe^{*}))$, where $(-)^{*}$ is the map induced by the involution $g\mapsto g^{-1}$.
@@
    Similarly, we define $T(RGe,RHf):=T(R[G\times H])(e\otimes f^{*})$ where we identify $R[G\times H]$ with $RG\otimes_{R}RH$ in the natural way. In the case where $e=f$, $T(RGe,RGe)$ has a ring structure induced by the tensor product $\otimes_{RGe}$, written $\cdot_{G}$. We denote by $T^{\Delta}(RGe,RHf)$ the subgroup of $T(RGe,RHf)$ the subgroup generated by all trivial source $(RGe,RHf)$-bimodules which have twisted diagonal vertices. Taking bimodule duals induces a group isomorphism $T(RGe,RHf)\to T(RHf,RGe)$, which has inverse again induced by taking duals. Denote by $O(T(RGe,RGe))$ the subgroup of $T(RGe,RGe)^{\times}$ consisting of all virtual $p$-permutation bimodules $\gamma$ whose inverse is its dual, i.e. $\gamma\cdot_{G}\gamma^{*}=[RGe]=\gamma^{*}\cdot_{G}\gamma$.
    Finally, observe that any $M\in T(RGe,RHf)$ induces a group homomorphism $M\cdot_{H}-:T(RHf)\to T(RGe)$. If $M\in T(RGe,RHf)^{\times}$, it is an isomorphism.
@@
    The following definition is due to Boltje and Xu in [BoXu07]. A $p$-permutation equivalence between $RGe$ and $RHf$ is an element $\gamma\in T^{\Delta}(RGe,RHf)$ satisfying
@@
Let $A$ be a sum of blocks of $RG$ and $B$ a sum of blocks of $RH$. A splendid Rickard equivalence for $A$ and $B$ is a bounded chain complex $C$ of $p$-permutation $(A,B)$-bimodules which satisfy the following conditions:
@@
    $C\otimes_{B}C^{*}\simeq A$ in $K^{b}({}_{A}\mathbf{mod}_{A})$ and $C^{*}\otimes_{A}C\simeq B$ in $K^{b}({}_{B}\mathbf{mod}_{B})$, where $A$ and $B$ are considered as chain complexes concentrated in degree 0.
A splendid Rickard equivalence $C$ of $(A,B)$-bimodules induces an equivalence of categories $C\otimes_{B}-:K^{b}({}_{B}\textbf{triv})\to K^{b}({}_{A}\textbf{triv})$.
@@
Since $B$ is symmetric and the terms of $C$ are projective as right $B$-modules, we have an isomorphism $C\otimes_{B}C^{*}\cong\mathrm{End}_{B^{op}}(C)$ of chain complexes of $(A,A)$-bimodules. First, suppose $\mathrm{End}_{B^{op}}(C)$ is contractible. The complex $\mathrm{End}_{A\otimes_{k}B^{op}}(C)$ is obtained from the complex $\mathrm{End}_{B^{op}}(C)$ by taking $A$-fixed points. Since taking $A$-fixed points is an additive functor, $\mathrm{End}_{A\otimes_{R}B^{op}}(C)$ is also contractible. Taking homology we obtain
@@
If $A$ and $B$ are blocks of $RG$ and $RH$, then a splendid Rickard equivalence for $A$ and $B$ is an indecomposable object in the homotopy category of chain complexes of $(A,B)$-bimodules.
@@
Suppose that $A$ is a block of $RG$ and $B$ is a block of $RH$. Let $C$ be a splendid Rickard equivalence for $A$ and $B$. Then $C$ is indecomposable in $K^{b}({}_{A}\mathbf{mod}_{B})$. In particular, $C$ has, up to isomorphism, a unique indecomposable noncontractible direct summand which is a splendid Rickard equivalence for $A$ and $B$.
@@
Historically, the notion of a “splendid” (SPLit ENDomorphism two-sided tilting complex Induced from Diagonal subgroups) Rickard equivalence is due to Rickard, and was additionally communicated by Broué. In Rickard’s paper [Ri96] such complexes are referred to as “splendid equivalences,” and the nomenclature “Rickard complex” was commonly used prior to refer to two-sided tilting complexes (see [Br93] for example, in which Rickard’s original definition also appears) due to his landmark paper [R91]. Rickard’s original definition assumed that $G$ and $H$ shared a common Sylow $p$-subgroup, and that the vertices of the modules in a splendid Rickard equivalence are contained inside the diagonal subgroup generated by the common Sylow. We remove that assumption here, instead asserting only that the modules are twisted diagonal.
@@
Assume that $A$ and $B$ are blocks of $RG$ and $RH$, respectively. Let $C$ be a bounded chain complex of $p$-permutation $(A,B)$-bimodules which satisfies condition 5.2(a). If $C\otimes_{B}C^{*}\simeq A$ in $K^{b}({}_{A}\mathbf{mod}_{A})$, then $C$ is a splendid Rickard equivalence.
@@
An indecomposable $(A,B)$-bimodule $M$ with a twisted diagonal vertex $X$ is projective as a left $A$-module and a right $B$-module. Indeed, by the Mackey Formula and the Krull-Schmidt theorem, every indecomposable direct summand of $\mathrm{Res}_{G\times 1}^{G\times H}(M)$ is projective relative to $(G\times 1)\cap{}^{(g,h)}X$ for some $(g,h)\in G\times H$. Since $^{(g,h)}X$ is also twisted diagonal, this intersection is trivial. This shows that $M$ is projective as a left $A$-module. A similar argument shows that $M$ is projective as a right $B$-module. Hence, the result follows from [Li181, 4.14.16]. ∎
@@
[BoXu07, Theorem 1.5] Let $C$ be a splendid Rickard equivalence between $A$ and $B$. Then $\gamma:=\Lambda(C)$ is a $p$-permutation equivalence between $RGe$ and $RHf$.
@@
Under Rickard’s original definition, if $C$ is a splendid Rickard equivalence for $A$ and $B$, for any $Q\leq P$ and blocks $e_{Q}$ and $f_{Q}$ corresponding to the blocks of $kC_{G}(Q)$ and $kC_{H}(Q)$ determined by $A$ and $B$, the chain complex $e_{Q}C(\Delta Q)f_{Q}$ is a splendid Rickard equivalence for $kC_{G}(Q)e_{Q}$ and $kC_{H}(Q)f_{Q}$. Such equivalences are regarded as “local” equivalences, equivalences between blocks of local centralizer subgroups.
@@ 6. Extended tensor products and tensor products of $p$-permutation bimodules
where $h\in H$ is chosen such that $(g,h)\in X$ and $(h,k)\in Y$. Note this is independent of choice of $h$. This construction was first used in [Bo10b]. We denote this extended tensor product by $M\underset{RH}{\overset{X,Y}{\otimes}}N\in{}_{R[X\ast Y]}\mathbf{mod}$, and obtain a bifunctor $\underset{RH}{\overset{X,Y}{\otimes}}-:{}_{RX}\mathbf{mod}\times{}_{RY}\mathbf{mod}\to{}_{R[X\ast Y]}\mathbf{mod}$.
Some quick calculations determine that this construction is associative in the following way: if $L$ is a finite group, $Z\leq K\times L$, and $P\in{}_{RZ}\mathbf{mod}$, then
@@
defines a natural isomorphism of $R[X\ast Y\ast Z]$-modules which is natural in $M,N$ and $K$.
Moreover, for $g\in G$, $h\in H$, $k\in K$, we have a natural isomorphism
@@
[Bo10b] Let $X\leq G\times H$, $Y\leq H\times K$, $M\in{}_{RX}\mathbf{mod}$, $N\in{}_{RY}\mathbf{mod}$. There is a natural isomorphism
@@
[BoPe20, Lemma 7.2] Suppose $(K,\mathcal{O},k)$ is a $p$-modular system. Let $X\leq G\times H$ and $Y\leq H\times K$.
@@
This construction originally appeared in [BoDa12]. For a fixed twisted diagonal $p$-subgroup $\Delta(P,\sigma,R)\leq G\times K$, denote by $\Gamma_{H}(P,\sigma,R)$ the set of triples $(\phi,Q,\psi)$ where $Q\leq H$ and $\psi:R\xrightarrow{\sim}Q$ and $\phi:Q\xrightarrow{\sim}P$ are isomorphisms with $\phi\circ\psi=\sigma$. The group $H$ acts on $\Gamma_{H}(P,\sigma,R)$ by
$^{h}(\phi,Q,\psi):=(\phi c_{h}^{-1},{}^{h}Q,c_{h}\psi).$
@@
(see [BoDa12, 3.1]) which is natural in $M$ and $N$. The following theorem characterizes the tensor product on the right-hand side for a particular class of $p$-permutation modules.
@@
[BoDa12, Theorem 3.3] Let $\Delta(P,\sigma,R)\in s_{p}(G\times K)$ be a twisted diagonal $p$-subgroup, let $\Gamma=\Gamma_{H}(P,\sigma,R)$ be as above, and let $\widetilde{\Gamma}\subseteq\Gamma$ be a set of representatives of the $H$-orbits of $\Gamma$. Furthermore, let $M\in{}_{\mathcal{O}G}\mathbf{triv}_{\mathcal{O}H}$ and $N\in{}_{\mathcal{O}H}\mathbf{triv}_{\mathcal{O}K}$ be $p$-permutation bimodules whose indecomposable direct summands have twisted diagonal vertices. Then the direct sum of the homomorphisms $\Phi_{(\phi,Q,\psi)}$ for $(\phi,Q,\psi)\in\widetilde{\Gamma}$ yields an isomorphism
@@
[BoPe20, Corollary 7.4] Let $M\in{}_{\mathcal{O}G}\mathbf{triv}_{\mathcal{O}H},N\in{}_{\mathcal{O}H}\mathbf{triv}_{\mathcal{O}K}$, $\Delta(P,\sigma,R)\in s_{p}(G\times K)$, and $\widetilde{\Gamma}\subseteq\Gamma=\Gamma_{H}(P,\sigma,R)$ be as before.
@@
    For the induced action of $N_{G\times K}(\Delta(P,\sigma,R))$ on the $H$-orbits $\phi,Q,\psi]_{H}$ of $\Gamma$, one has
@@
[BoPe20, Theorem 7.5] Let $(P,e)\in\mathcal{BP}(kG)$ and $(R,d)\in\mathcal{BP}(kK)$. Suppose $\sigma:R\xrightarrow{\sim}P$ is an isomorphism and let $C_{G}(P)\leq S\leq N_{G}(P,e)$ and $C_{K}(R)\leq T\leq N_{K}(R,d)$ be intermediate subgroups. Further, let $\Omega:=\Omega_{H}((P,e),\sigma,(R,d))$ denote the set of triples $(\phi,(Q,f),\psi)$, where $(Q,f)\in\mathcal{BP}(kH)$ and $\psi:R\xrightarrow{\sim}Q$ and $\phi:Q\xrightarrow{\sim}P$ are isomorphisms such that $\sigma=\phi\circ\psi$. Finally, let $M\in{}_{\mathcal{O}G}\mathbf{triv}_{\mathcal{O}H}$ and $N\in{}_{\mathcal{O}H}\mathbf{triv}_{\mathcal{O}K}$ be $p$-permutation modules all of whose indecomposable direct summands have twisted diagonal vertices.